# Old Quantum Mechanics by Bohr and Sommerfeld from a Modern Perspective

### v2

@@ Abstract.
Primary 81-01, 81-03. Secondary 81CDedicated to the 100th anniversary of the birth of wave mechanics [52].
@@ 1. Introduction
Among the key sources on the so-called “Old Quantum Mechanics” of Bohr and Sommerfeld are the classic publications [1], [8], [20], [30], [31], [40], [44], [49], [56], [57], [66], and several educational videos [9]. In this paper, we aim to explore the following topics from a mathematical perspective:
Topics: (i) The Bohr model: circular orbits of electrons in hydrogen-like atoms; derivation of the Bohr formula (Nobel Prize in Physics, 1922 [7]). (ii) Wilson and Sommerfeld: quantization rules for multi-dimensional periodic systems via classical action; Sommerfeld’s relativistic formula for elliptical orbits. (iii) Elementary evaluation of Sommerfeld-type integrals. (iv) Additional examples and a resolution of the “Sommerfeld puzzle”. (v) Appendix A: Vector calculus tools for uniform circular motion. (vi) Appendix B: Instability of the hydrogen atom in classical physics due to the electron’s fall into the center, as predicted by Rutherford’s model. (vii) Appendix C: Independent evaluation of Sommerfeld-type integrals using parameter differentiation.
Brief history: The fine structure of hydrogen atom spectral lines was discovered by Albert A. Michelson in 1887 [42, 43]. When his ether-wind experiments failed, he turned to spectroscopy and found that the prominent line of the Balmer series was actually a doublet [6, 49]. The electron was discovered by J. J. Thomson in 1897 [63]. Rutherford proposed the planetary model of the atom in 1911. In 1916, Arnold Sommerfeld extended the quantization rules of the ‘old’ quantum theory to the relativistic hydrogen atom [57] (see also [23, 29, 44, 49]). An exact solution was achieved only in 1928 by C. G. Darwin [11] and W. Gordon [22] after the discovery of the Dirac equation [14, 16] – astonishingly, the new result precisely matched the ‘old’ Sommerfeld formula – the so-called “Sommerfeld Puzzle” [6]!
@@ 2. Bohr’s Atomic Model
Mendeleev’s Periodic Table and the Bohr model are two significant developments in understanding the structure of elements and atoms — Mendeleev’s table, created in 1869, organized elements by increasing atomic weight and recurring chemical properties [56, pp. 2–3]. The Bohr model, proposed in 1913 [8], provided a model for the structure of the atom, depicting electrons orbiting the nucleus in fixed energy levels. The simplified way of understanding the electronic structure of atoms is directly related to the arrangement of elements in the periodic table.
@@ 3. Wilson and Sommerfeld Quantization Rules in Wave Mechanics
Topics to review: The spherical harmonics [35, 46, 64], Schrödinger equation [13, 35, 51], relativistic Schrödinger and Dirac equations [3, 13, 51, 62]; the spinor spherical harmonical harmonics [61, 62, 64] and separation of variables for the Dirac equation in a central field, semiclassical approximation [46, 51].
@@
In the development of quantum theory, the Bohr–(Wilson)–Sommerfeld quantization rule served as an original “bridge” between classical and quantum mechanics (for historical details, see [39], [40], [44], [49], [57], [58], and [66]). Nowadays, we use the Schrödinger [55] and Dirac [15] wave equations for the corresponding Kepler problems. How did Schrödinger derive his celebrated equation and subsequently apply it to the hydrogen atom? According to his own testimony [52, 53, 54], de Broglie’s seminal work on a wave theory of matter (1923–24) [12] and Einstein’s work on ideal Bose gases (1924–25) laid the foundation for the discovery of wave mechanics (see also [3] and [41]).
@@
This approximation refines the Bohr–Sommerfeld quantization rule within wave mechanics. The WKB method, which provides approximate solutions to wave equations, leads to a quantization condition similar to the Bohr–Sommerfeld rule, with a slight modification involving a phase shift. During separation of variables, the quantization of angular momentum and spin is exact, since the concept of spin is already embedded into the structure of the corresponding wave equation [33].
@@
The corresponding wave functions, their relations to Airy functions [17], quantization rules, and further technical details are discussed in [46, 51], and elsewhere. (We recommend the reader to review sections 19, pp. 235–251, and 28, pp. 178–188, on the semiclassical approximation of [46, 51], respectively.)
@@
For all Coulomb problems under consideration, we utilize a generic integral, originally evaluated by Sommerfeld [58] using complex analysis: If
@@
with . (See Section 4 and Appendix 9 for two independent elementary evaluations of this integral.)
@@
(see Figure 3 for the original version222Schrödinger’s notebooks are reproduced in the Archive for the History of Quantum Physics (AHQP); for more details, see [27, 41]. ) and applies Langer’s transformation to define the effective momentum:
@@
However, Schrödinger’s relativistic approach fails to accurately describe the fine structure of hydrogen-like atoms (hydrogen, ionized helium, doubly-ionized lithium, etc.). For instance, the total fine-structure splitting at is a factor of too large compared to Sommerfeld’s prediction, which agrees with experimental observations.333The maximum spread in the fine-structure levels occurs for and with total angular momentum and in Eqs. (3.19) and (3.24), respectively [13, 51]. The ratio of these spreads is:
@@
This prediction agrees with experimental data on fine-structure splittings for the hydrogen-like systems.
@@
For a full analytical solution, including the non-relativistic limit, see Refs. [19, 46, 61, 62] (based on the Nikiforov–Uvarov method), as well as standard texts [13, 51].
@@ 4. Evaluation of the Sommerfeld-Type Integrals
For all problems under consideration, we utilize the generic integral (3.7)–(3.8), originally evaluated by Sommerfeld using complex integration [58] (see Figure 4). We now discuss an elementary evaluation of this integral [3]. Integrating by parts on the left-hand side of (3.8), one obtains:
@@ 5. Further Examples and Resolution of “Sommerfeld’s Puzzle”
As is well known, Bohr introduced his semiclassical quantization rules for hydrogen-like atoms based on classical circular motion, but Sommerfeld extended these ideas to relativistic elliptical orbits. Measurements of the fine structure by Paschen were interpreted as experimental tests of the special theory of relativity [6, 29]. The exact solution was obtained, for the first time, by C. G. Darwin [11] and W. Gordon [22], only after the discovery of the Dirac equation [14, 16] – the new answer was precisely the ‘old’ Sommerfeld formula (3.22)!
@@
Indeed, the generic (WKB-based) rule (3.9) is also valid for the exact energy levels obtained via the Nikiforov–Uvarov approach [19, 60] for all Coulomb problems under consideration. Other examples include quantum harmonic oscillators, and systems with Kratzer and Pöschl–Teller potentials [19]. (See also [34] for an extension of Schrödinger’s coherent states [53].)
@@
Note. It can be summarized that only after the “Two Quantum Revolutions” the ambiguity of quantization of the Kepler problem in the ‘old quantum mechanics’ [23, 48] was resolved, when the spherical symmetry had been taking into account by separation of variables and the corresponding radial equations were derived, without any assumptions, subject to further WKB approximation under the Langer transformation. Schrödinger seems to have done it first in 1925 in his unpublished notes but he used equation (3.14), which corresponds to the spin zero particle (see [3], Figure 3 for his original notebook, and [33] for further details on the concept of spin and the wave equations). As was later attested, due to discrepancy with experimental data, he never published this work (see also [3], Appendix D, for his letter to Weyl).
@@ 6. Conclusion
Traditional textbooks [13, 35, 51] do not discuss the derivation of the Sommerfeld fine structure formula in the semiclassical approximation, and now the reader can understand why! Indeed, de facto, there are three different levels of complexity, as in the historical “Three Quantum Revolutions”: – Elementary one, in the Bohr model of atom. – Introductory quantum mechanics for the non-relativistic hydrogen atom. – And finally, relativistic quantum theory, the Dirac equation, for the fine structure formula. The WKB method can be applied only after a thorough study of basic properties of the Dirac equation, including the construction of spinor spherical harmonics [61, 62, 64] followed by a non-trivial separation of variables in spherical coordinates. Our notes, while not perfect, may help the reader fill in these gaps. The resolution of the “Sommerfeld Puzzle” is also of interest [60]. Needless to say, practical use of computer algebra [2, 19] will definitely help students in learning quantum physics, for instance, when performing tedious calculations.
@@ 8. Appendix B: Instability of a Hydrogen Atom in Classical Physics
As is known [9, 49, 56]a rotating electron in Rutherford’s planetary model must fall into the nucleus, the spiral-in, according to the laws of classical mechanics and electrodynamics. The total instantaneous power emitted over all solid angles is given by the well-known result of Larmor [26]:

### v3

@@ Abstract.
Primary 81-01, 81-03. Secondary 81CPrediction is very difficult, especially if it’s about the future!
Niels Bohr
If you want to be a physicist, you must do three things - first, study mathematics, second, study more mathematics, and third, do the same.
Arnold Sommerfeld
@@ 1. Introduction
Among the key sources on the so-called “Old Quantum Mechanics” of Bohr and Sommerfeld are the classic publications [1, 12, 23, 26, 39, 40, 49, 55, 60, 68, 69, 72, 80] and several educational videos [13]. In this paper, we aim to explore the following from a mathematical perspective:
Topics: (i) The Bohr model: circular orbits of electrons in hydrogen-like atoms; derivation of the Bohr formula (Nobel Prize in Physics, 1922 [10]). (ii) Wilson and Sommerfeld: quantization rules for multi-dimensional periodic systems via classical action; Sommerfeld’s relativistic formula for elliptical orbits. (iii) Elementary evaluation of Sommerfeld-type integrals. (iv) Additional examples and a resolution of the “Sommerfeld puzzle”. (v) Appendix A: Vector calculus tools for uniform circular motion. (vi) Appendix B: Instability of the hydrogen atom in classical physics due to the electron’s fall into the center, as predicted by Rutherford’s model. (vii) Appendix C: Independent evaluation of Sommerfeld-type integrals using parameter differentiation. (viii) Appendix D: Letter from Schrödinger to Sommerfeld dated January 29, 1926.
Brief history: The fine structure of hydrogen atom spectral lines was discovered by Albert A. Michelson in 1887 [53, 54]. When his ether-wind experiments failed, he turned to spectroscopy and found that the prominent line of the Balmer series was actually a doublet [8, 38, 60]. The electron was discovered by J. J. Thomson in 1897 [76]. Rutherford proposed the planetary model of the atom in 1911. In 1916, Arnold Sommerfeld extended the quantization rules of the ‘old’ quantum theory to the relativistic hydrogen atom [69] (see also [31, 38, 55, 60, 70]). An exact solution was achieved only in 1928 by C. G. Darwin [15] and W. Gordon [30] after the discovery of the Dirac equation [18, 20] – astonishingly, the new result precisely matched the ‘old’ Sommerfeld formula – the so-called “Sommerfeld Puzzle” [8] and [23, pp. 426–429]!
@@
Our goal: These notes are intended as a supplement to traditional textbooks [17, 44, 62] and our recent article [4], offering our own explanations, insights, historical context, and expanded discussion on specific topics. They can be useful when teaching and learning quantum physics and for an honor project at any level. It is motivated by an introductory course in mathematics of quantum mechanics which one of the authors (SKS) has been teaching at Arizona State University for more than two decades [24, 75].
@@ 2. Bohr’s Atomic Model
Mendeleev’s Periodic Table and the Bohr model are two significant developments in understanding the structure of elements and atoms — Mendeleev’s table, created in 1869, organized elements by increasing atomic weight and recurring chemical properties [68, pp. 2–3]. The Bohr model, proposed in 1913 [12], provided a model for the structure of the atom, depicting electrons orbiting the nucleus in fixed energy levels. The simplified way of understanding the electronic structure of atoms is directly related to the arrangement of elements in the periodic table.
@@ 3. Wilson and Sommerfeld Quantization Rules in Wave Mechanics
Topics to review: Kepler problems [8], [27, pp. 146–148], [28, pp. 92–102, pp. 466–477, pp. 481–482], [70, pp. 84–90, pp. 109–119, pp. 251–258]; spherical harmonics [24, 44, 57, 77], Schrödinger equation [17, 44, 52, 62], relativistic Schrödinger and Dirac equations [2, 4, 6, 17, 62, 75]; the spinor spherical harmonical harmonics [2, 6, 74, 75, 77] and separation of variables for the Dirac equation in a central field, semiclassical approximation [5, 29, 57, 62].
@@
In the development of quantum theory, the Bohr–(Wilson)–Sommerfeld quantization rule served as an original “bridge” between classical and quantum mechanics (for historical details, see [48], [49], [55], [60], [69], [70], and [80]). Nowadays, we use the Schrödinger [67] and Dirac [19] wave equations for the corresponding Kepler problems. How did Schrödinger derive his celebrated equation and subsequently apply it to the hydrogen atom? According to his own testimony [63, 65, 66] and [51, 030† pp. 141–143],222In this letter to Einstein on November 3, 1925, he writes: A few days ago, I read with great interest the ingenious theses of Louis de Broglie, which I finally got hold of … . de Broglie’s seminal work on a wave theory of matter (1923–24) [16] and Einstein’s work on ideal Bose gases (1924–25) laid the foundation for the discovery of wave mechanics (see also [4] and [50]).
@@
This approximation refines the Bohr–Sommerfeld quantization rule within wave mechanics. The WKB method, which provides approximate solutions to wave equations, leads to a quantization condition similar to the Bohr–Sommerfeld rule, with a slight modification involving a phase shift. During separation of variables in spherical coordinates, the quantization of angular momentum and spin is exact, since the concept of spin is already embedded into the structure of the corresponding wave equation [42].
@@
More to review: The WKB wave functions, their relations to Airy functions [21], quantization rules, and further technical details are discussed in [57, 62], and elsewhere. We recommend the reader to review sections 19, pp. 235–251, and 28, pp. 178–188, on the semiclassical approximation of [57, 62], respectively; as well as, chapter 9 of [21] on Airy functions.
@@
For all Coulomb problems under consideration, we utilize a generic integral, originally evaluated by Sommerfeld [70, pp. 611–612] using complex analysis: If
@@
with (see also [28, pp. 468–470]). In Section 4 and Appendix 10 we present two independent elementary evaluations of this integral.
@@
(see Figure 3 for the original version333Schrödinger’s notebooks are reproduced in the Archive for the History of Quantum Physics (AHQP); for more details, see [35, 50].) and applies Langer’s transformation to define the effective momentum:
@@
This prediction agrees with experimental data on fine-structure splittings for the hydrogen-like systems. On the contrary, Schrödinger’s relativistic approach fails to accurately describe the fine structure of hydrogen-like atoms (hydrogen, ionized helium, doubly-ionized lithium, etc.). For instance, the total fine-structure splitting at is a factor of too large compared to Sommerfeld’s prediction, which agrees with experimental observations. The maximum spread in the fine-structure levels occurs for and with total angular momentum and in Eqs. (3.19) and (3.24), respectively [17, 62]. The ratio of these spreads is:
When one gets
Note. With the help of Mathematica we derived the next two terms in (3.24) as follows [3, 70]:
and
@@
For a full analytical solution, including the non-relativistic limit, see Refs. [24, 57, 74, 75] (based on the Nikiforov–Uvarov method), as well as standard texts [2, 6, 17, 62].
@@ 4. Evaluation of the Sommerfeld-Type Integrals
Teaching mathematics in the United States is in favor of “real world problems” — this is the one! For all problems under consideration, we utilize the generic integral (3.7)–(3.8), originally evaluated by Sommerfeld using complex integration [70, pp. 611–612] (see Figure 4). We now discuss an elementary evaluation of this integral [4]. Integrating by parts on the left-hand side of (3.8), one obtains:
@@ 5. Further Examples and Resolution of “Sommerfeld’s Puzzle”
As is well known, Bohr introduced his semiclassical quantization rules for hydrogen-like atoms based on classical circular motion, but Sommerfeld extended these ideas to relativistic elliptical orbits. Measurements of the fine structure by Paschen were interpreted as experimental tests of the special theory of relativity [8, 38, 70, 72]. The exact solution was obtained, for the first time, by C. G. Darwin [15] and W. Gordon [30], only after the discovery of the Dirac equation [18, 20] – the new answer was precisely the ‘old’ Sommerfeld formula (3.22)!
@@
For exact solutions, in the corresponding generalized equation of the hypergeometric type [57],
we may choose
Then
must be a linear function [57]. When one can complete the square and obtain
We may choose
As a result, for all Sommerfeld-type potentials, by the Nikiforov–Uvarov quantization rule [57],
one obtains
as a generic equation for the exact energy levels with (It is worth noting that Sommerfeld had obtained a similar relation in special cases [71].)
The puzzle resolution:
Theorem 1.
Indeed, the generic (WKB-based) rule (3.9) is also valid for the exact energy levels (5.7) obtained via the Nikiforov–Uvarov approach [24, 73] for all Coulomb problems under consideration. Other examples include quantum harmonic oscillators, and systems with Kratzer and Pöschl–Teller potentials [24]. (See also [43] for an extension of Schrödinger’s coherent states [65].)
@@
Methodological note: It should be clear by now that only after the “Two Quantum Revolutions” the ambiguity of quantization of the Kepler problem in the ‘old quantum mechanics’ [31, 59] can be resolved, when the spherical symmetry is taking into account by separation of variables and the corresponding radial equations are derived exactly, without any assumptions, subject to further WKB approximation under the Langer correction.
Schrödinger seems to have been the first to go this route in 1925 (or was close to that but did avoid a mistake?) in his unpublished notes but he used equation (3.14), which corresponds to the spin zero particle (see [4], Figure 3 for his original notebook, and [42] for further details on the concept of spin and the wave equations). As was later attested, due to discrepancy with experimental data, he never published this work (see also [4, Appendix D] for his letter to Weyl).
6. A Mistake that Schrödinger Never Made
Interestingly, in Figure 3, the bottom left shows the ‘old’ Bohr–Sommerfeld quantization rule with what is now known as the Langer correction [5], namely, but with instead of on the right hand side, as it is supposed to be in the WKB approximation (3.6) (in Schrödinger’s notation you may want to interchange as one has to write nowadays; see also footnote12 on p. 18 below). As we now understand, this would have resulted in the wrong spectrum for his relativistic equation. But Schrödinger never made this mistake and, instead, outlined, “…(although having used just written words)…” [4, Appendiz D, p. 100], the exact solution by the Laplace method in terms of a contour integral [4, Appendix C]!
Schrödinger knew the Sommerfeld-type integral very-well. For example, it is explicitly cited in his letter to Sommerfeld on January 29, 1926: [50, p. 462] and [51, 041† pp. 170–172]: “… Finally, I still wish to add that the discovery of the whole connection [between the wave equation and the quantization of hydrogen atom], goes back to your beautiful quantization method for evaluating the radial quantum integral. It was the characteristic which suddenly shone out from the exponents and like a Holy Grail.”
In this letter Schrödinger reported, for the first time, the success of his non-relativistic theory in the case of quantum oscillator, rotator, and the hydrogen atom (Kepler problem) before the forthcoming publications. He also formulated a program for future research. For the benefits of the reader, the entire letter is translated from German to English in Appendix D.
Timetable: The exact dates of Schrödinger’s fundamental discoveries in his first publications [63, 64] are not recorded [36, 37, 79] and [50, pp. 459–465]. But one can take into account his letter to Einstein [51, 030† pp. 141–143], Bloch’s recollection of two colloquiums in Zürich [9], letters to Wein [51, 037† pp. 162–165] from Arosa and to Sommerfeld from Zürich [51, 041† pp. 170–172]. This gives us an estimate from the beginning of November 1925 to the end of January 1926.
At the same time, in January 1926, Bohr looked back on the development of the ‘old’ theory in a letter to his friend, the Swedish physicist Carl Oseen [38, p. 85](quoted from [11, p. 73]): “At the present stage of the development of the quantum theory we can hardly say whether it was good or bad luck that the properties of the Kepler motion could be brought into such simple connection with the hydrogen spectrum, as was believed possible at one time. If this connection had merely had that asymptotic character which one might expect from the correspondence principle, then we should not have been tempted to apply mechanics as crudely as we believed possible for some time. On the other hand, it was just these mechanical considerations that were helpful in building up the analysis of the optical phenomena which gradually led to quantum mechanics.” It was hard to predict, indeed!
@@ 7. Conclusion
Traditional textbooks [2, 6, 17, 44, 52, 62] do not discuss the derivation of the Sommerfeld fine structure formula in the semiclassical approximation, and now the reader can understand why! Indeed, de facto, there are three different levels of complexity, as in the historical “Three Quantum Revolutions” in the development of quantum physics: – Elementary one [27, 28], in the Bohr model of atom. – Introductory quantum mechanics [44, 52], for the non-relativistic hydrogen atom. – And finally, relativistic quantum theory [2, 6, 17, 62], the Dirac equation, for the fine structure formula. The WKB method can be applied only after a thorough study of basic properties of the Dirac equation, including the construction of spinor spherical harmonics [2, 6, 74, 75, 77] followed by a non-trivial separation of variables in spherical coordinates. Obviously, all this represents a “pedagogical challenge”. Our notes, while not perfect, may help the reader fill in these gaps. The resolution of the “Sommerfeld Puzzle” is also of interest [73]. Needless to say, practical use of computer algebra [3, 24] will definitely help students in learning quantum physics, for instance, when performing tedious calculations.
@@ 9. Appendix B: Instability of a Hydrogen Atom in Classical Physics
As is known [13, 60, 68], a rotating electron in Rutherford’s planetary model must fall into the nucleus, the spiral-in, according to the laws of classical mechanics and electrodynamics. The total instantaneous power emitted over all solid angles is given by the well-known result of Larmor [34]:
@@ 10. Appendix C: An Independent Evaluation of the Sommerfeld-type Integrals
11. Appendix D: A Letter from Schrödinger to Sommerfeld
Schrödinger to Sommerfeld [51, 041† pp. 170–172]
Zurich, January 29, 1926555This letter is also printed in Arnold Sommerfeld, Scientific Correspondence, Vol. 2, pp. 236–238 [In German]:
Arnold Sommerfeld, Wissenschaftlicher Briefwechsel, Band 2, S. 236–238
Most Honored Professor,
It has been so long since I have not let you hear from me,666Schrödinger’s last (extant) letter [51, 025† pp. 132–135] to Sommerfeld was written in July 1925. that I must now write quickly, so that you don’t write to me even earlier, after you have seen my quantum work, which I sent to Privy Councillor Wien for the Annalen777This was Schrödinger’s first communication on wave mechanics [63], which he had sent a few days earlier to one of the editors of Annalen der Physik., with the request that I will show it to you in advance. Of course, I am more curious not about anyone’s judgment but than about yours, whether you share the very high hopes I have for the derivation of quantum laws from a Hamiltonian principle?888Cf. [27, 28] – See also Sommerfeld’s reply [51, 042† pp. 173–175].
Since then, I’ve incorporated a few more mechanical problems into the new conception. As far as my mathematics goes, everything unfolds in the most beautiful way - and yet is not a copy of the old quantum rules, but rather differs in characteristic points.999According to Schrödinger’s former colleague Peter Paul Ewald [25, p. 385], Courant and Hilbert’s seminal 1924 work on methods of mathematical physics “enabled physicists to grasp the spirit of a unified mathematical method in an especially important field, marked by keywords: eigenvalues, eigenfunctions. These problems… arise in physics through oscillations of all kinds.”
The linear oscillator must be treated with the same analytical tools as the Kepler problem (the square of the abscissa must be introduced as an independent variable in the oscillation equation obtained for the function ).101010The “Planck oscillator” is treated in Schrödinger’s second paper [64] as the first example. Once again, the strange case arises that an equation that cannot be integrated by ordinary quadratures can be integrated, precisely for the eigenvalues, even by elementary functions - a proof of how kindly nature is to facilitate our understanding. The eigenvalues (energy levels) are: i.e., so-called half-integer quantization.111111Cf. [64]. Half-integer quantum numbers had also been introduced in Heisenberg’s early attempts to describe the anomalous Zeeman effect. Even if the quantum differences remain unchanged here, I still see this as a good omen, because is the arithmetic mean of and
The rotator (dumbbell model) in three dimensions (i.e., with the two variables ) is quite simple; the eigenfunctions are ordinary spherical harmonics, and the eigenvalues (energy levels) are This characteristic value comes from the differential equation of the spherical harmonics. I find it very pleasing, not precisely because of the present case, but because it gives hope that with further development, instead of 121212As Sommerfeld noted in [70, pp. 333, 476], the appearance of instead of in the Landé -factor “suggests that not one state but two adjacent quantum states and are physically relevant.” will also be obtained at those points where it is needed (your intensity formulas and the formulas for the anomalous Zeeman splitting).131313In the introduction to his supplement on wave mechanics [71, p. 2], Sommerfeld pointed out that these difficulties of the old quantum theory had now found their natural explanation through the new wave mechanics.
In the case of the rotator, on the contrary, one would first have to consider whether the explanation of the band spectra would be harmful. But I hardly think so. The effect, as one easily sees, is a very small difference in the linear term in the positive and negative branches, and if I remember correctly, it actually exists anyway (or something like that; I only calculated all of this yesterday and haven’t been able to check yet).
For the free motion of a point mass, we obtain that any energy value is possible if the point mass is located in infinite space. If it is located in a box, which is to be regarded as a boundary condition for the -function, one obtains discrete energy levels, approximately the same as those obtained by quantizing the zigzag motion. The eigenfunctions for the free mass point, if one calculates relativistically, correspond to de Broglie phase waves;141414As we learn from his letter to Einstein [51, 030† pp. 141–143], Schrödinger was inspired to this work by the “brilliant theses” (1925) of Louis de Broglie on the phase waves of electrons, which he had first encountered in early November 1925. for the mass point in the box, they are standing eigen-oscillations of the box volume with the dispersion law of de Broglie phase waves.
The next most important task, apart from the calculation of important special cases, such as the Stark effect, the Zeeman effect, and relativistic Keplerian motion, appears to be the establishment of a rule for intensity and polarization, which must replace the correspondence principle.151515See also the remarks in his letters to Lorentz [51, 055† pp. 203–205] and Wentzel [51, 068† pp. 226–228]. Since I believe that the -function truly describes the processes in the atom that cause the emission of light, it must provide information about them. One must investigate the intensity beats between two simultaneously excited eigenvibrations and their (the beats’) spatial distribution. The rich mathematical theory (orthogonality of eigenfunctions, significance of eigenvalues as extreme values of the Hamiltonian integral, etc.) will surely lead to simple theorems. I am concerned about the relativistic Kepler problem.161616Schrödinger, as shown in his notes (Figure 3), had first proposed a relativistic wave equation before arriving at the well-known non-relativistic one [4]. See also his letters to Sommerfeld [51, 044† pp. 178–184] and Lorentz [51, 076† pp. 252–261]; together with historical studies [36, 37, 50]. I am not sure whether it is true what I say in the paper, that nuclear co-motion is so essential in the new way of treating it. Even less would I maintain my observation that it might already be the case with the old method of treatment, and would ask you to delete this remark if you consider it nonsense.171717Sommerfeld addressed this point in his next letter [51, 042† pp. 173–175].
But for me, of course, the help can only come from the nuclear motion, otherwise you get half-quanta – contradicting experience. And these half-quanta stem precisely from that of the spherical functions, which, on the other hand, is so welcome.
Finally, I would like to say that the discovery of the entire connection, even if it is not externally apparent, goes back to your beautiful complex integration method for evaluating the radial quantum integral. It was the characteristic and familiar which suddenly shone out from the exponents and like a Holy Grail.181818Cf. [70, p. 611–612].
I hope that you, most esteemed Professor, and all your family are well. With best and most devoted greetings from house to house,
I remain always
Yours faithfully, E. Schrödinger
P. S. Innsbruck has not yet been officially decided.191919Schrödinger had also received an offer from Innsbruck (cf. also his letter to Sommerfeld [51, 025† pp. 132–135] ), which he declined in mid-March (cf. his letter [51, 052† pp. 197–200] of March 17, 1926, to Thirring). But I think I’ll stay here. It is mainly Schweidler’s departure to Vienna202020The previous Innsbruck chair, Egon von Schweidler (1873–1948), had just been appointed to the I. Physics Institute in Vienna. His successor in Innsbruck, after the chair was downgraded to a lectureship (cf. [51, 044† pp. 178–184]), became Schrödinger’s friend Arthur March (1891–1957). that decides for me. Herzfeld wrote to me after speaking with you (I don’t know if it was directly your opinion) that we should try to support Smekal.212121Adolf Smekal (1895–1959) did not receive an extraordinary professorship in Vienna until 1927, and in 1928 he accepted a call to Halle. Herzfeld was at that time preparing to accept a professorship at Johns Hopkins University in Baltimore. I think it will be difficult because March is named que es loco and has been supplying for quite some time. But Thirring wants to do it anyway and wrote to me at the time, asking me to inform him after an official decision has been made. It would certainly be more appropriate from a factual perspective.
Please don’t assume my decision is final during the discussion. It would be unpleasant for both ministries. And on the other hand, I quite appreciate the delay because, after much effort, I’ll finally get a little something, namely a new blackboard in the lecture hall and, hopefully, a little more funding for the seminar library.